# Agent Control Protocol ACP v1.19: Admission Control for Agent Actions Technical Specification and Reference Implementation

### 25 Mar 2026 v5

@@ 16.1  Adversary Capabilities
We consider a probabilistic adversary $\mathcal{A}$ interacting with the system through agent requests. The adversary is assumed to have the following capabilities:
  • Request control: $\mathcal{A}$ can generate arbitrary sequences of requests on behalf of one or more agents, including high-frequency and adversarial patterns.
  • Adaptive behavior: $\mathcal{A}$ can adapt future requests based on past responses (approval/denial outcomes), enabling iterative probing strategies.
  • Multi-agent coordination: $\mathcal{A}$ may coordinate multiple agents to attempt distributed evasion, including alternating identities or cross-agent patterns.
  • Input manipulation: $\mathcal{A}$ can choose all request fields (capability, resource, context) within the bounds of the protocol specification.
@@ 16.3  Security Properties
For a given request and system state, the evaluation function produces a unique and reproducible decision outcome. This prevents adversarial exploitation of non-determinism. The RiskDeterminism invariant in the TLA+ model formally specifies this property.
@@
Policy enforcement is applied independently per agent. Actions by one agent do not affect the enforcement state of another. This property bounds the impact of a compromised or malicious agent to its own request history.
@@ 16.4  Informal Security Argument
Because ACP separates stateless evaluation from state management and enforces a strict execution contract, every request results in a deterministic state transition that cannot be bypassed. The append-only nature of the state backend ensures that historical information accumulates monotonically: cooldown activation follows deterministic threshold conditions applied to accumulated state; repeated probing eventually triggers enforced denial periods.
@@ 16.5  Limitations
  • Post-quantum signature migration (ACP-SIGN-2.0 HYBRID, Ed25519 $\to$ ML-DSA-65) is specified and stubbed; full implementation is deferred to v1.20.
ACP is therefore best understood as a control-layer mechanism that assumes a trusted execution and storage environment, while providing strong guarantees on decision consistency and temporal behavior under adversarial inputs.
17  Related Work
17.1  Access Control and Policy Enforcement
17.2  Runtime Policy Systems and Agent Frameworks
Recent work has begun to address security governance for LLM-based agentic systems. Syros et al. [15] propose SAGA, a security architecture that provides user-controlled access policies and cryptographic mediation for agent-to-agent tool use and inter-agent communication. ACP complements this line of work by focusing on deterministic, auditable admission control at the individual request level: rather than governing inter-agent communication channels, ACP enforces stateful risk evaluation and cooldown constraints on each action request, with an externally verifiable execution contract and formal invariants.
17.3  Auditability and Tamper-Evident Logs
17.4  Formal Verification of Protocols
ACP incorporates formal methods in a targeted manner: the TLA+ model (tla/ACP.tla) mechanically checks three critical invariants (fail-closed admission, append-only ledger, risk determinism) using TLC. Rather than attempting full-system verification, ACP adopts a pragmatic approach in which formally verified invariants are instantiated as concrete sequence test vectors and validated at runtime by the ACR-1.0 compliance runner. This creates a three-layer verifiability chain: formal $\to$ data $\to$ runtime.
17.5  Summary
Property RBAC ABAC OPA Cedar MCP/A2A ACP
Stateful decisions $\times$ $\times$ $\times$ $\times$ $\times$
Deterministic evaluation $\times$
Auditability partial partial partial partial $\times$
Runtime enforcement partial
Temporal behavior $\times$ $\times$ $\times$ $\times$ $\times$
External verifiability $\times$ $\times$ $\times$ $\times$ $\times$
Execution contract $\times$ $\times$ $\times$ $\times$ $\times$
Agent-native $\times$ $\times$ $\times$ $\times$ partial
Table 20: Comparison of ACP with existing access control and agent governance approaches. ✓ = fully supported; $\times$ = not supported; partial = partially addressed.
@@ 18  Conclusion
ACP is not the first attempt to control autonomous agents. It is the first attempt to do so through a formal technical specification with precise state models, demonstrable security properties, and verifiable conformance requirements. The difference between a best-practices policy and a formal protocol is exactly that: behaviors are defined, failures have specific error codes, and conformance can be verified.
@@
The v1.19 specification is complete. A full Go reference implementation (23 packages, L1–L4), 73 signed + 65 unsigned RISK-2.0 conformance test vectors, 5 stateful sequence test vectors (ACR-1.0 compliance runner), a TLC-runnable TLA+ formal model (3 invariants, 0 violations), an ACP-SIGN-2.0 HYBRID mode stub with Ed25519 and ML-DSA-65 migration path, an executable multi-organization interoperability demo (examples/multi-org-demo/), a payment-agent killer demo (examples/payment-agent/), and an extended OpenAPI 3.1.0 specification (18 endpoints) are publicly available at https://github.com/chelof100/acp-framework-en.
Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19219776
@@
ACP demonstrates that admission control for autonomous agents can be both computationally efficient and operationally enforceable—provided that decision logic and state management are cleanly separated. The protocol is compute-cheap but state-sensitive: the LedgerQuerier abstraction keeps the evaluation function at $\sim$820 ns while allowing the state backend to be replaced, optimized, or scaled independently. This is not an implementation detail—it is the architectural property that makes ACP deployable in production without sacrificing governance correctness.
@@
info@traslaia.com | https://agentcontrolprotocol.xyz
@@ Appendix A Glossary
Appendix B Formal Verification (ACP-RISK-2.0 — TLC-Runnable)
The tla/ACP.tla module is a complete TLC-runnable TLA+ model of the ACP-RISK-2.0 admission control engine. It checks three invariants over a bounded state space (2 agents $\times$ 4 capabilities $\times$ 3 resources, ledger depth $\leq 5$):
@@
Interpretation. Safety encodes fail-closed admission: every APPROVED decision had $\mathit{RS}\leq 39$. LedgerAppendOnly encodes tamper-evidence in a state-based view: every ledger entry carries a valid decision and a non-negative risk score. LedgerAppendOnlyTemporal is the temporal statement: in every step, existing entries are preserved and the ledger never shrinks. RiskDeterminism is the central invariant of ACP-RISK-2.0: identical (capability, resource) pairs always produce identical risk scores, enabling third-party recalculation from a signed policy snapshot.
Scope of the formal model. The TLA+ model intentionally focuses on the deterministic core of the protocol—admission evaluation, ledger append-only property, and risk score determinism—and does not cover all operational aspects of ACP. Behavioral properties such as cooldown state transitions, multi-step delegation sequences, and anomaly counter evolution are verified empirically through the ACR-1.0 sequence compliance runner and its five stateful test vectors. Extension of the formal model to cover these behaviors is left for future work.

### 30 Apr 2026 v10

@@ 19.1  Adversary Capabilities
We formalize the adversary as a triple $\mathcal{A}=(K,S,B)$ where:
  • $K\in\{\textit{black-box},\ \textit{formula-aware},\ \textit{full-state}\}$ is the knowledge level. A black-box adversary observes only the outcome of each request (APPROVED/ESCALATED/DENIED). A formula-aware adversary additionally knows the published risk formula ($\mathrm{RS}=\mathrm{capBase}+F_{\mathrm{res}}+F_{\mathrm{ctx}}+F_{\mathrm{hist}}+F_{\mathrm{anom}}$) and all weight values defined in ACP-RISK-3.0. A full-state adversary further has real-time read access to the entire ledger state (pattern counts, denial history, cooldown timestamps).
  • $S\in\{\textit{threshold-hugging},\ \textit{evasion},\ \textit{flood},\ \textit{collusion}\}$ is the strategy class. Threshold-hugging: craft requests with $\mathrm{RS}$ just below enforced boundaries. Evasion: systematically drive all features toward zero to achieve permanent APPROVED. Flood: saturate the system with high-volume requests. Collusion: coordinate multiple agents to distribute request load across isolation boundaries.
  • $B\in\mathbb{N}$ is the request budget available to the adversary.
The evaluation experiments instantiate specific $(K,S,B)$ triples. Table 43 maps each experiment to its adversary class and reports the enforcement outcome.
Table 43: Adversary taxonomy: mapping of ACP experiments to $(K,S,B)$ adversary classes.
Exp Adversary Class $K$ $S$ $B$ Outcome
1 Cooldown Evasion black-box threshold-hugging 500 contained
2 Multi-Agent Flood black-box flood+collusion 1000 contained
3 Backend Stress black-box flood 10K degraded perf
4 Token Replay black-box evasion 300 contained
5–8 State Mixing formula-aware evasion fixed in RISK-3.0
9 Deviation Collapse upstream drift BAR detects
10 Knowledge-Aware Evasion formula-aware evasion (RS=0) 200 BAR detects
11 Threshold Sensitivity monotonic
12 Multi-Tool IPI Chain formula-aware flood 8 stateful persistence
13 Coordination Window black-box round-robin/burst $5N$ linear bound ($O(N)$)
[…]
Five findings emerge from this taxonomy:
  1. 1.
    Per-decision enforcement suffices against black-box adversaries. Experiments 1–4 confirm that stateful admission control contains all black-box strategies within the request budget.
  2. 2.
    Formula-aware adversaries break per-decision enforcement but are detected by BAR. Experiment 10 shows that a formula-aware adversary can achieve $\mathrm{RS}=0$ on every request, causing every decision to be APPROVED — yet BAR-Monitor fires a TREND alert at batch 2, before the regime reaches full collapse.
  3. 3.
    Deviation collapse (Exp 9) is an environmental threat, not adversarial, but BAR-Monitor handles both classes with the same mechanism.
  4. 4.
    Stateful ledger creates persistent enforcement consequences (Exp 12). An IPI flood that triggers cooldown also activates the 24-hour anomaly window ($F_{\text{anom}}$ Rule 2), elevating borderline capabilities from ESCALATED to DENIED for the entire anomaly window after cooldown expires. A stateless engine would return to pre-attack RS values immediately.
  5. 5.
    Full-state adversaries are outside the v1.0 threat model. An adversary with real-time ledger access could predict and exploit threshold transitions. This class is not addressed in the present specification and is acknowledged in §13.8 as a target for ACP-D-1.0 (decentralized governance).
  6. 6.
    Coordinated multi-agent activity scales linearly (Exp 13). Per-agent behavioral isolation ensures that $N$ agents acting in concert can execute at most $2N$ approved operations before the first DENIED decision—a coordination window that grows linearly with $N$, with zero deviation observed. Superlinear amplification through coordination is structurally prevented by the combination of agent-scoped pattern keys and the evaluate-then-mutate execution contract.
@@ 19.3  Security Properties
ACP determinism is defined with respect to the evaluation context: given identical canonical request inputs, system state (ledger, pattern counters, cooldown state), policy definitions, and evaluation order, the protocol produces a consistent and reproducible admission decision. This form of determinism explicitly includes state as part of the input to the decision function; it does not imply independence from time or state evolution. This scoping prevents adversarial exploitation of non-determinism across evaluators. The RiskDeterminism and AnomalyDeterminism invariants in the TLA+ model formally specify this property across all reachable states.
@@
Policy enforcement is applied independently per agent. Actions by one agent do not affect the enforcement state of another. This property bounds the impact of a compromised or malicious agent to its own request history. Note that isolation holds across agents but not across capability contexts within the same agent: anomaly signals from F_anom Rules 1 and 2 aggregate at the agent level, enabling cross-context interference for agents holding multiple capability tokens under the same keypair (see §13.8, “State-Mixing and Context Fragmentation”).
@@ 19.4  Informal Security Argument
Because ACP separates stateless evaluation from state management and enforces a strict execution contract, every request results in a deterministic state transition that cannot be bypassed. When the state backend enforces append-only semantics, historical information accumulates monotonically: cooldown activation follows deterministic threshold conditions applied to accumulated state; repeated probing eventually triggers enforced denial periods.
@@ 19.5  Limitations
  • Distributed deployments: ACP’s determinism guarantee holds per evaluator instance. In deployments with replicated or eventually-consistent state backends (e.g., Redis Cluster, Postgres streaming replicas), two evaluators may observe divergent ledger states and reach different admission decisions for the same agent at the same point in time. ACP does not specify a distributed consensus protocol; deployments requiring strict cross-instance consistency must enforce strong consistency at the storage layer.
@@
  • Prompt-layer attacks and LLM-specific benchmarks. ACP operates at the execution layer and does not prevent indirect prompt injection (IPI) [22] at the prompt layer. An IPI attack that causes an agent to emit a low-risk request (e.g., acp:cap:data.read/Public, $\mathrm{RS}=0$) will be admitted by ACP, because the resulting tool call is individually valid. ACP’s enforcement boundary is the tool call, not the instruction that produced it. Standard LLM agent benchmarks—AgentDojo [8], InjecAgent [51], and AgentContract-Bench—evaluate prompt-layer manipulation resistance, which is a complementary but orthogonal property to execution-layer admission control. Prompt-layer defenses reduce the IPI attack surface; ACP enforces structural constraints over the action sequence regardless of how the request was generated. The two layers are composable: a prompt-layer filter applied before the tool call reaches the ACP admission boundary reduces the frequency of injected high-risk requests, while ACP provides the structural guarantee that high-risk requests are denied regardless of their origin. ACP’s defense against formula-aware evasion (Experiment 10, §17.8) addresses a distinct threat class: an adversary who drives $\mathrm{RS}=0$ across all requests—which BAR-Monitor detects via $\Delta\mathrm{BAR}$, not per-decision enforcement.
  • Post-quantum signature migration (ACP-SIGN-2.0 HYBRID, Ed25519 + ML-DSA-65) is implemented via the Cloudflare CIRCL library (pkg/sign2/). ML-DSA-65 signing latency is approximately 100–130 µs (vs. 25 µs for Ed25519), remaining well below state-backend latency; see §12.2.
ACP is therefore best understood as a control-layer mechanism whose guarantees are conditioned on the integrity of the state backend and the execution environment. When these components operate correctly, ACP provides strong guarantees on decision consistency and temporal behavior under adversarial inputs. When the state backend is compromised (e.g., history deleted by a privileged operator), the governance guarantees collapse; this is a declared residual risk, not a design oversight. Deployments in high-assurance environments should pair ACP with tamper-evident storage—a cryptographically chained log structure that makes deletion and modification detectable (motivating a future ACP-LEDGER-1.0 specification).
@@ 20  Conclusion
ACP is not the only attempt to control autonomous agents. It advances this goal through a formal technical specification with precise state models, empirically validated security properties, and verifiable conformance requirements. The difference between a best-practices policy and a formal protocol is exactly that: behaviors are defined, failures have specific error codes, and conformance can be verified.
@@
The v1.25 specification is complete. A full Go reference implementation (23 packages, L1–L4), 73 signed + 65 unsigned RISK-2.0 conformance test vectors, 5 stateful sequence test vectors (ACR-1.0 compliance runner), a TLC-runnable TLA+ formal model (base: 3 invariants; extended v1.25: 11 invariants + 4 temporal properties, 5,684,342 states generated, 3,147,864 distinct, graph depth 15, 0 violations; two-agent safety at LB=11: 4,648,235,723 states generated, 4,294,930,695 distinct, 11 invariants, 0 violations, 10.5 h), a real ML-DSA-65 (Dilithium mode3) hybrid implementation via Cloudflare CIRCL (pkg/sign2/, SignHybridFull + VerifyHybrid), an executable multi-organization interoperability demo (examples/multi-org-demo/), a payment-agent demo (examples/payment-agent/), an extended OpenAPI 3.1.0 specification (18 endpoints), ten experiments covering performance benchmarks across backend configurations (Experiments 1–4) and adversarial evaluation: token replay (Experiment 5), stateless vs. stateful comparison (Experiment 6), state-mixing vulnerability characterization (Experiment 7), context-scoped anomaly enforcement under ACP-RISK-3.0 (Experiment 8), deviation collapse and restoration (Experiment 9), and knowledge-aware adversarial evasion (Experiment 10) are publicly available at https://github.com/chelof100/acp-framework-en [13].
Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19672575
@@
Cross-agent coordination remains a necessary extension for deployments involving adversarial multi-identity scenarios, particularly in high-integrity systems; a concrete mitigation path is described in §13.8 (Distributed Attack Surface).
ACP demonstrates that admission control for autonomous agents can be both computationally efficient and operationally enforceable—provided that decision logic and state management are cleanly separated. The protocol is compute-cheap but state-sensitive: the LedgerQuerier abstraction keeps the evaluation function at 739–832 ns (p50; Table 12) while allowing the state backend to be replaced, optimized, or scaled independently. This is not an implementation detail—it is the architectural property that makes ACP deployable in production without sacrificing governance correctness.
A system that always produces admissible actions does not demonstrate effective governance; it may instead be operating in a regime where admissibility is no longer meaningfully evaluated. The contribution of this work is to make that condition detectable, measurable, and formally specifiable.
@@
info@traslaia.com | https://agentcontrolprotocol.xyz [12]
@@ Appendix A Glossary
Appendix B Prohibited Behaviors
ACP defines 12 behaviors that no compliant implementation may exhibit. Exhibiting any of the following disqualifies conformance at all levels.
Code Prohibited behavior
PROHIB-001 Approving a request when any evaluation component fails
PROHIB-002 Reusing an already-consumed Execution Token
PROHIB-005 Allowing state transition from revoked
PROHIB-007 Modifying or deleting Audit Ledger events
PROHIB-008 Silencing ledger corruption detection
PROHIB-009 Ignoring max_depth in delegation chains
PROHIB-011 Approving requests from agents with autonomy_level 0
Table 45: ACP prohibited behaviors (ACP-CONF-1.2). Exhibiting any disqualifies conformance at all levels.
Appendix C Formal Verification (ACP-RISK-2.0 — TLC-Runnable)
The tla/ACP.tla module is a complete TLC-runnable TLA+ model of the ACP-RISK-2.0 admission control engine. The full TLA+ model and TLC configuration files are available in the project repository at https://github.com/chelof100/acp-framework-en [13] (tla/ directory). It checks three invariants over a bounded state space (2 agents $\times$ 4 capabilities $\times$ 3 resources, ledger depth $\leq 5$). Liveness properties in the extended model depend on two explicit assumptions: monotonic time progression (time does not freeze indefinitely) and weak fairness in request evaluation (an enabled request is eventually processed). These assumptions are declared in the Spec formula via WF_vars(Tick) and WF_vars(EvaluateRequest) respectively; they are standard for temporal verification and are stated here to make the verification scope transparent.
Model scope.
The TLA+ specification abstracts $F_{\text{ctx}}$ (context factors) and $F_{\text{hist}}$ (history-based signals) from the risk computation. Verified properties hold over this abstracted model; their applicability to the full ACP-RISK-2.0 formula is supported empirically via the ACR-1.0 conformance runner and the experiments in §13.7, but is not formally model-checked. Table 46 summarizes what is verified, what is empirically tested, and what falls outside the current formal scope.
Table 46: Verification coverage of ACP-RISK-2.0 components. ✓ = covered; $\times$ = not covered in that layer.
Component / Signal TLA+ verified Empirically tested Notes
Cooldown enforcement Temporal property (CooldownExpires)
Flood detection (Rule 1) Extended model (ACP_Extended.tla)
Denial escalation (Rule 2) Verified via execution traces
Pattern anomaly (Rule 3) Context-isolated (patternKey)
Context factors ($F_{\text{ctx}}$) $\times$ $\times$ Abstracted; future formal work
History factors ($F_{\text{hist}}$) $\times$ Validated via ACR-1.0 traces
State-mixing behavior $\times$ Measured in Experiment 7
@@ Appendix B Prohibited Behaviors
Interpretation. Safety encodes fail-closed admission [40]: every APPROVED decision had $\mathit{RS}\leq 39$. LedgerAppendOnly encodes tamper-evidence in a state-based view: every ledger entry carries a valid decision and a non-negative risk score. LedgerAppendOnlyTemporal is the temporal statement: in every step, existing entries are preserved and the ledger never shrinks. RiskDeterminism is the central invariant of ACP-RISK-2.0: identical (capability, resource) pairs always produce identical risk scores, enabling third-party recalculation from a signed policy snapshot. In the base model, ComputeRisk is a pure function of its inputs, so RiskDeterminism holds by construction. Its non-trivial content arises in the extended model, where AnomalyDeterminism extends the guarantee to include per-agent state: identical (capability, resource, pattern_at_eval) triples always produce the same risk score, establishing that RS is a deterministic policy function of the execution trace rather than a statistical estimate.
Extended model (v1.20): cooldown, delegation integrity, and Fanom flood detection. The tla/ACP_Extended.tla module extends ACP.tla with per-agent cooldown temporal state, denial accumulation, a static delegation chain, and—added in Sprint J2—per-agent request pattern accumulation ($F_{\text{anom}}$ flood detection). It introduces six additional invariants and three temporal properties beyond ACP.tla:
  • CooldownEnforced (safety): active cooldown forces DENIED for any request by the affected agent, regardless of risk score.
  • CooldownImpliesThreshold (safety): cooldown only exists after the denial threshold (COOLDOWN_TRIGGER) has been reached—cooldown is never spurious.
  • DelegationIntegrity (safety): no consecutive self-delegation in the chain (chain[i] $\neq$ chain[i+1], ACP-DCMA-1.1 §3).
  • AnomalyDeterminism (safety, Sprint J2): same (capability, resource, pattern_at_eval) triple always produces the same risk score. Extends RiskDeterminism to include per-agent state: $\text{RS}=f(\text{static inputs}+\text{request history count})$. RS is a deterministic policy function, not a statistical model.
  • FloodEnforced (safety, Sprint J2): when the per-agent request count meets or exceeds FLOOD_THRESHOLD at evaluation time, the decision is DENIED—the $F_{\text{anom}}$ flood override is never bypassed.
  • CooldownExpires (liveness): active cooldown eventually expires, conditioned on the expiry time being within the model’s time horizon (cooldown_until $\leq$ MAX_TIME). Requires weak fairness on time advance (WF_vars(Tick)) in Spec.
  • EventuallyRejected (liveness, Sprint J2): an abusive agent not already in cooldown will eventually receive a DENIED decision via the flood override. Requires per-agent weak fairness on EvaluateRequest and LEDGER_BOUND >> FLOOD_THRESHOLD.
  • NoInfiniteExecutionUnderAbuse (liveness, Sprint J2, strong form): no execution trace exists in which an agent exhibiting sustained abusive behavior (request count $\geq$ FLOOD_THRESHOLD) continues to be admitted indefinitely. Formally: $\Box(\mathit{Abusive}(a)\Rightarrow\neg\Diamond(\Box\mathit{Accepted}(a)))$. Holds in the bounded model via EventuallyRejected; holds in the unbounded specification by structural induction on FloodActive monotonicity (pattern_count is non-decreasing; once $\geq$ FLOOD_THRESHOLD, all subsequent decisions are DENIED).
Extended TLC result (Sprint J2 configuration). Running java -jar tla2tools.jar -deadlock -config ACP_Extended.cfg ACP_Extended.tla with $|\mathit{Agents}|=1$, $|\mathit{Capabilities}|=3$, $|\mathit{Resources}|=2$, LEDGER_BOUND=7=7, FLOOD_THRESHOLD=4=4, FANOM_BONUS=25=25, COOLDOWN_TRIGGER=3=3, COOLDOWN_WINDOW=3=3, MAX_TIME=7=7:
Model checking completed. No error has been found.
5,684,342 states generated, 3,147,864 distinct states found, 0 states left on queue.
The depth of the complete state graph search is 15.
Finished in 47min 30s (TLC2 v2.16, Java 1.8, single worker, 7241MB heap).
The model captures agent behavior under repeated request patterns, incorporating stateful anomaly detection and cooldown enforcement. All eleven invariants and all four temporal properties hold with zero violations across the complete bounded state space, under the stated assumptions of monotonic time progression, weak fairness in request evaluation, and consistent atomic state updates. The graph depth of 15 confirms that the model exercises multiple execution paths and state transitions prior to enforcement—abusive behavior can persist across multiple steps before triggering rejection, confirming that enforcement is not immediate and that liveness properties are not satisfied trivially.
D4 liveness result. In particular, the model verifies a liveness property ensuring bounded execution under sustained abuse. No execution trace was found in which an agent exceeding the flood threshold could continue to be admitted indefinitely. This result provides evidence that, under the stated assumptions and within the bounded model, the protocol enforces eventual containment of abusive behavior rather than allowing unbounded execution under repeated high-risk patterns.
Single-agent configuration is used for $F_{\text{anom}}$ liveness verification; per-agent enforcement is independent by design (pattern_count[a] and last_decision[a] are per-agent maps), so the result generalizes structurally to $N$ agents.
Two-agent safety check (J2c). The model was further evaluated under a two-agent configuration (Agents={‘‘A1’’,‘‘A2’’}, FLOOD_THRESHOLD=3, LEDGER_BOUND=5, MAX_TIME=5) to assess potential interference effects between agents. Due to the exponential growth of the state space under multi-agent interaction, reduced bounds were used for this configuration. All nine safety invariants hold across all 1,650,024 distinct states (3,011,690 states generated, graph depth 11, 1 h 08 min) with zero violations, confirming per-agent enforcement isolation: pattern_count[A1], cooldown state, and admission decisions for A1 do not affect A2’s evaluation state, and vice versa.
Temporal properties were not enforced in this configuration, as liveness guarantees depend on continued request generation and sufficient time progression. Under reduced bounds, executions may terminate without additional requests, leading to spurious liveness violations that do not reflect protocol behavior under sustained interaction. The observed behavior highlights an important modeling consideration: temporal guarantees in ACP assume ongoing interaction with the system. In bounded executions where request generation ceases, the model may reach terminal states in which no further enforcement transitions occur; these traces do not represent sustained adversarial behavior and are therefore outside the intended scope of the liveness guarantees. Accordingly, liveness properties are evaluated under the single-agent configuration (FLOOD_THRESHOLD=4, MAX_TIME=7), where bounds are calibrated to allow sufficient interaction depth.
Two-agent safety at scale (J2d). To strengthen the isolation claim, a further two-agent run was conducted with bounds calibrated to match the single-agent liveness configuration: Agents={‘‘A1’’,‘‘A2’’}, FLOOD_THRESHOLD=4, LEDGER_BOUND=6, MAX_TIME=7, with 6 parallel workers and 12 GB heap (TLC2 v2.16, Java 26). The run completed in 4 h 13 min, generating 50,927,772 states and finding 27,648,560 distinct states at graph depth 14. All eleven safety invariants hold with zero violations across the complete bounded multi-agent state space, confirming that per-agent enforcement isolation extends to the full invariant set under the liveness-calibrated bounds.
TLC reported a temporal property violation at the final liveness check, but was unable to generate a counter-example trace: the heap was exhausted (OutOfMemoryError) while constructing the lasso over 165 million accumulated state-space entries. Diagnosis reveals a model-bounding artifact rather than a protocol defect. With two agents sharing a ledger of capacity LEDGER_BOUND=6=6, the combined request volume fills the ledger before either agent reaches FLOOD_THRESHOLD=4=4, disabling EvaluateRequest for both agents and preventing EventuallyRejected from firing. The necessary condition for liveness verification in an $N$-agent configuration is:
$\texttt{LEDGER\_BOUND}\;\geq\;\texttt{FLOOD\_THRESHOLD}\times N+\delta,$
[…]
Two-agent safety at extended scale (LB=11). To directly address the scaling constraint identified above, a further run was conducted at the required bound: Agents={‘‘A1’’,‘‘A2’’}, LEDGER_BOUND=11, FLOOD_THRESHOLD=4, MAX_TIME=7, with 15 parallel workers (auto-scaled by TLC) and 24 GB heap on an Intel Core i9-13900HX with 32 GB RAM (TLC2 v2.16, Java 21, Windows 11). The run explored 4,294,930,695 distinct states over 10.5 hours, generating 4,648,235,723 total states, with all eleven safety invariants reporting zero violations throughout the entire exploration. The run was halted by disk space exhaustion (584 GB of checkpoint data written to disk), not by any property violation. This result confirms that all per-agent enforcement invariants hold robustly under the correct multi-agent ledger bound, providing safety verification at $756\times$ the scale of the single-agent configuration and $155\times$ the scale of the LEDGER_BOUND=6 run. Liveness verification at this scale remains open: the complete state space at LEDGER_BOUND=11 substantially exceeds current hardware limits for lasso-trace computation.
Exploration collapse: a formal analog of deviation collapse. The bounding artifact above exhibits a structural parallel to the runtime phenomenon studied in Section 17. In deviation collapse, an upstream sanitizer eliminates the risk signals that would cause enforcement to fire: the engine remains active and syntactically correct, but its boundary is never exercised. In the multi-agent TLC configuration, bounded shared resources eliminate the state conditions under which liveness properties can be exercised [37, 1]: the model checker remains sound and the invariants hold, but the traces required to validate temporal guarantees cannot be reached within the bound. Unlike classical state-space reductions—such as partial-order methods, which preserve liveness properties by construction [5]— resource-induced bounds may eliminate entire classes of trajectories from the reachable state space. We refer to this as resource-induced liveness obstruction: a structural limitation that arises not from a defect in the protocol or the model, but from the interaction between bounded shared resources and multi-agent concurrency.
Both cases instantiate the same structural pattern: when the conditions necessary for enforcement to fire disappear—whether from the runtime pipeline or from the formal model—validation loses meaning independently of whether the mechanism itself is correct. Put concretely: systems appear compliant because no invalid actions occur; models appear correct because no violations are reachable. In both cases, the decision boundary remains defined and the enforcement logic remains intact, but the capacity to exercise failure conditions has been eliminated. This observation motivates a modeling requirement analogous to failure condition preservation (Section 17): formal liveness verification of multi-agent systems requires that resource bounds be sufficient to generate the behavioral conditions under which liveness properties are exercised [50, 29]. Addressing this in the general case—including per-agent ledger allocation, explicit fairness constraints, and scalable model checking strategies—is identified as future work.
ACP thus provides two distinct guarantees: safety in all executions (invariants hold regardless of scheduling or agent count, verified across 4,294,930,695 distinct states in the two-agent configuration at the correct ledger bound LEDGER_BOUND=11) and liveness under sustained interaction (eventual enforcement holds when agents continue to generate requests within the time horizon, verified for the single-agent configuration where bounds are calibrated to admit the relevant traces).
Design notes on the extended model. The base risk score ComputeRisk(cap, res) is preserved as a pure function; the extended function ComputeRiskWithAnom(cap, res, pat_count) adds FANOM_BONUS when pat_count $\geq$ FLOOD_THRESHOLD, modeling ACP-RISK-2.0 $F_{\text{anom}}$ Rule 3 (rapid-fire anomaly accumulation). FloodActive(a) additionally overrides the decision to DENIED regardless of the resulting risk score, ensuring the flood enforcement cannot be bypassed by capability/resource combinations with inherently low base scores. Denial counts increment only for RS-based DENIED decisions—not for cooldown-forced or flood-forced denials—consistent with ACP-RISK-2.0 §4 (AddDenial fires on real denials only). The delegation chain is a static 2-hop sequence; dynamic chain mutation is outside the scope of the bounded model (reserved for a future version). The CooldownExpires and EventuallyRejected liveness properties are conditioned on the relevant state being reachable within the time horizon; the unconditional forms hold in the unbounded specification.
Framing. Sprint J2 extends formal coverage from cooldown enforcement to flood-based $F_{\text{anom}}$ enforcement: the model now verifies not only that individual denials are correct (Safety, CooldownEnforced) but also that sustained abusive behavior is blocked within the bounded model under the stated assumptions (NoInfiniteExecutionUnderAbuse). The AnomalyDeterminism invariant formalizes the key correctness property of ACP’s stateful risk engine: the admission decision is fully reproducible given the same capability, resource, and per-agent request history, enabling independent audit without proprietary infrastructure. The phrase “formally verified” is deliberately avoided; the correct claim is: model checking of selected safety and liveness properties under a bounded state model, under the stated assumptions of monotonic time progression, weak fairness in request processing, and consistent atomic state updates.
Model simplifications. The TLA+ model uses ComputeRiskWithAnom(cap, res, pat_count) in place of the full ACP-RISK-2.0 formula, which additionally includes $F_{\text{ctx}}$ (context factors) and $F_{\text{hist}}$ (history factors). $F_{\text{anom}}$ is modeled as a frequency-based accumulation plus a direct override (FloodActive); semantic payload analysis is out of scope. This abstraction is intentional: it captures a fundamental class of abuse patterns (sustained high-frequency interaction from the same agent) while enabling tractable model checking. The key contribution is not the detection signal itself, but the verifiable enforcement guarantee—the formal proof that, given a deterministic accumulation function with defined thresholds, the protocol enforces correct admission decisions under stateful conditions, produces auditable execution traces, and allows independent reproduction by third parties. This distinguishes ACP from conventional rate limiters, which lack formal specification, execution contracts, and external verifiability. pattern_count is monotone within a bounded trace (no per-window decay), approximating the time-windowed behavior of ACP-RISK-2.0 $F_{\text{anom}}$ Rule 3. The model verifies properties for representative parameter values (FLOOD_THRESHOLD=4=4, FANOM_BONUS=25=25); complete generalization to arbitrary parameter values is not claimed. Full determinism under $F_{\text{hist}}$ is validated empirically via the ACR-1.0 compliance runner and the adversarial evaluation in Section 13.7. These simplifications are deliberate: the model is intended to verify structural safety and liveness properties, not to replicate the full engine.
Limitations of the formal evaluation. The model checking is performed over a bounded model with a finite time horizon and a limited number of agents. As such:
  • The results do not constitute a proof of correctness for unbounded executions.
  • Liveness guarantees rely on assumptions of monotonic time progression and weak fairness in request evaluation.
  • A supplementary two-agent safety check (J2c) confirms per-agent enforcement isolation across 1,650,024 distinct states with 0 violations; temporal liveness properties are verified only in the single-agent configuration with calibrated bounds. Full multi-agent contention under adversarial coordination is not modeled.
  • The state space is constrained by bounded parameters (time horizon, ledger size), which may not capture all possible long-horizon behaviors.
These constraints define the scope of the guarantees provided by the model and should be considered when interpreting the results.